# Knowledge Independence Breeds Disruption but Limits Recognition

### v1

@@ Acknowledgments
The seminal work of Uzzi et al. (?) demonstrated that papers built on atypical combinations of prior work are more likely to be highly influential. This finding is based on a novelty index, which quantifies the degree of atypicality in the combinations of references at the journal level. More specifically, this index is calculated by comparing the observed co-occurrence of journal pairs in actual citation networks with their expected co-occurrence under randomized citation networks. This comparison allows for computing the $Z$-score of each possible journal pair cited by a given focal paper. A high and positive $Z$-score indicates a conventional combination of journal pairs, while a low and negative $Z$-score reflects a novel combination that deviates significantly from expected co-occurrence patterns. Following the way that novelty is computed in the SciSciNet  (?) dateset, we calculate the $10_{\rm th}$ percentile of the $Z$-score distribution of the journal pairs cited by the focal paper, and apply its negative value to represent the $90_{\rm th}$ percentile of novelty. This approach underscores the relatively top atypical combinations of references within the paper, capturing its degree of novelty.
@@
Lin et al.  (?) reported the rise of remote collaborations that span locations and time zones. These collaborations are less likely to produce disruptive works as they tend to focus more on technical tasks rather than conceptual ones. The original calculation of team distance relies on the latitude and longitude of the affiliations listed in a paper’s byline. Due to the absence of coordinate information in the Web of Science dataset, we instead use the following hierarchical categorization of the affiliations:
  • 00: Sub-organization / Department level;
  • $1$: Organization / University level;
  • $2$: City level;
  • $3$: State / Province level;
  • $4$: National level;
  • $5$: International level.
Using this hierarchy, the distance of any given team is categorized based on the highest level within the team. For instance, if there are three authors, two of which are in the same university, while the third is in another university situated at the same city, the distance will be categorized as $2$, implying that this is a city-level collaboration.
@@
Zeng et al.  (?) revealed a positive correlation between team freshness and scientific disruption, emphasizing the importance of new team dynamics in fostering innovative research. To quantify team freshness, we analyze the prior collaboration network of the authors involved in a given focal paper. In this network, the undirected edges represent prior collaborations between pairs of authors before coauthoring the focal paper. We classify team freshness into mutually exclusive categories based on the topological structure of the collaboration network, where $V$, $k$, and $n$ represent the node set, node degree, and network size, respectively. More specifically, the classification is as follows:
@@
  • 00: All authors have previously collaborated with each other, $\{\forall v_{i}\in V,\ k_{i}=n-1\}$;
  • $1$: All authors have at least one prior collaboration link, $\{\forall v_{i}\in V,1\leq\ k_{i}\leq n-1\}$;
@@
  • $2$: Some authors have at least one prior collaboration link, $\{\exists v_{i}\in V,\ k_{i}=0\}$;
  • $3$: No author has any prior collaboration link, $\{\forall v_{i}\in V,\ k_{i}=0\}$.
Propensity score matching (PSM)
Propensity score matching (PSM)  (?) has been extensively employed to estimate causal treatment effects in the Science of Science  (?, ?). It offers a robust approach to account for potential biases induced by confounding factors, particularly in scenarios where randomized experiments are impractical. In this study, we apply PSM to rigorously quantify the effect size of KI on scientific disruption. The procedure involves the following steps. We first categorize papers into the treated group and controlled group according to treatment size (KI). Subsequently, we estimate the propensity score for each paper by fitting a generalized linear model where the treatment size is the dependent variable. The model includes the following covariates: focal paper properties (impact $C_{5}$ and publication year), reference properties (novelty, reference count, average age, average impact $C_{5}$, and average disruption), and team properties (size, geographic distance, and collaboration freshness). For each paper $i$ in the controlled group, a counterpart from the treated group with a similar propensity score is matched. This matching ensures a balanced distribution of covariates across the two groups, effectively mimicking the conditions of a randomized controlled trial. After identifying $m$ matched pairs of papers, denoted as $M$, we calculate the Average Treatment Effect on the Treated as $\text{ATT}=\frac{1}{m}\sum_{i\in M}(Y^{1}_{i}-Y^{0}_{i})$, where $Y^{1}_{i}$ is the outcome (scientific disruption) for the treated paper and $Y^{0}_{i}$ is the outcome for the control paper. To enhance the robustness of PSM results, we perform multiple experiments for different pairs of treated and control groups. This produces a matrix of ATT values (Fig. 2E), providing a comprehensive and nuanced view of the effect size of KI on scientific disruption.
@@
To analyze the relationships between scientific disruption and various covariates, we employ normalized ordinary least squares (OLS) regression models. This approach allows us to evaluate whether KI has a dominant effect on disruption. The regression models include the following covariates: focal paper properties (impact $C_{5}$, discipline, and publication year), reference properties (novelty, reference count, average age, average impact $C_{5}$, and average disruption), as well as team properties (size, geographic distance, and collaboration freshness). All numerical variables are standardized by $Z$-score normalization to ensure comparability of effect coefficients across models. We start with a model that does not control for any covariates other than fixing discipline and publication year (S1). Then, we control for each of the following covariates separately (S2): focal paper’s impact ($C_{5}$), novelty, reference count, reference age, reference impact ($C_{5}$), reference disruption, team size, team distance, and team freshness. Finally, we consider a model that accounts for all those covariates (S3). The results are displayed in Table 1.
$D$ $=b_{0}+a_{0}\text{KI}+c_{0}F,$ (S1)
$D$ $=b_{0}+a_{0}\text{KI}+a_{i}C_{i}+c_{0}F,$ (S2)
$D$ $=b_{0}+a_{0}\text{KI}+\sum\nolimits_{i=1}^{9}a_{i}C_{i}+c_{0}F.$ (S3)
@@
Over the past decades, mediation analysis has seen rapid methodological development  (?), becoming particularly prevalent in biomedical and social sciences. This approach is often applied to data from randomized controlled trials, aiming to investigate whether and how the effect of a target variable $X$ on an outcome $Y$ operates through a mediating variable $M$. Here we utilize the model-based mediation framework, comprising two parametric regression models: Outcome Model (S4) formulates the relation between the outcome $Y$ and the target variable $X$, as well as the mediator $M$, with covariates $C$ controlled. On the other hand, the Mediator Model (S5) formulates the relation between the mediator $M$ and the target variable $X$, also controlling for covariates $C$. The indirect effect (ACME) is defined as the product of the coefficient $\beta_{1}$ of $X$ in model (S5) and the coefficient $\theta_{2}$ of $M$ in model (S4). This represents the portion of $X$’s effect on $Y$ mediated through $M$. The direct effect (ADE) is defined as the coefficient $\theta_{1}$ of $X$ in model (S4), capturing the direct relationship between $X$ and $Y$, independent of $M$. The total effect (TE) is defined as the sum of the indirect effect (ACME) and direct effect (ADE). In all conducted mediation analyses, we standardize the numerical variables by $Z$-score normalization to ensure comparability of effect coefficients across models, with discipline and publication year controlled.
$Y$ $=\theta_{0}+\theta_{1}X+\theta_{2}M+\theta_{3}C,$ (S4)
$M$ $=\beta_{0}+\beta_{1}X+\beta_{2}C.$ (S5)
@@
Figure S1: Distribution of reference list length. The distribution of reference count for 33,708,219 WOS articles follows a stretched exponential pattern, with a large number of papers containing relatively short reference lists.
@@
Figure S2: The association between disruption and KI is robust across different reference list lengths. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group ($<=10$), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
@@
Figure S3: The association between disruption and KI is robust across alternative measures. (A-D) Upper panel: from the perspective of percentile, all disruption measures (materials and methods) show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of $i$-type references compared to $j$-type references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
@@
Figure S4: The association between disruption and KI is robust when controlling for novelty. (A) The novelty index (materials and methods) within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on less novel combinations of references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for more novel combinations of references. This pattern remains consistent across time periods. (B-C) When controlling for focal papers’ novelty, both the percentile and positive ratio of disruption continue to increase with KI across all levels of novelty. Additionally, papers relying on more novel combinations of references tend to exhibit slightly higher disruption.
@@
Figure S5: The association between disruption and KI is robust when controlling for reference age. (A), We represent the reference age of a given focal paper by calculating the the average age of cited references. The reference age within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on relatively younger references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for older references. This pattern remains consistent across time periods. (B-C), When controlling for reference age, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference age.
Figure S6: The association between disruption and KI is robust when controlling for reference impact. (A), We represent the reference impact of a given focal paper by calculating the average $C_{5}$ of cited references. The reference impact within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite less impactful references. This pattern remains consistent across time periods. (B-C), When controlling for reference impact, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference impact. Additionally, papers citing less impactful references tend to exhibit slightly higher disruption.
@@
Figure S7: The association between disruption and KI is robust when controlling for reference disruption. (A) We represent the reference disruption of a given focal paper by calculating the average disruption of cited references. The reference disruption within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite more disruptive references. This pattern remains consistent across time periods. (B-C) When controlling for reference disruption, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference disruption. Additionally, papers citing more disruptive references tend to exhibit higher disruption.
@@
Figure S8: The association between disruption and KI is robust when controlling for team size. (A-B) When controlling for team size, both the percentile and positive ratio of disruption continue to increase with KI across different team sizes. Additionally, papers produced by smaller teams tend to exhibit slightly higher disruption.
@@
Figure S9: The association between disruption and KI is robust when controlling for team distance. (A-B) We categorize the team distance of a given focal paper by hierarchical categorization of the affiliations, where the value from 00 to $5$ refers to collaboration hierarchy from sub-organization level to international level (materials and methods). When controlling for team distance, both the percentile and positive ratio of disruption continue to increase with KI across different team distances. Additionally, papers produced by geographically closer teams tend to exhibit slightly higher disruption.
@@
Figure S10: The association between disruption and KI is robust when controlling for team freshness. (A-B) We classify the team freshness of a given focal paper by the topological structure of the prior collaboration network, where the value from 00 to $3$ refers to network topology from a full-connected network to an isolated network (materials and methods). When controlling for team freshness, both the percentile and positive ratio of disruption continue to increase with KI across different team freshnesses. Additionally, papers produced by fresher teams tend to exhibit slightly higher disruption.

### v2

@@ Methods
The seminal work of Uzzi et al.  [uzzi2013atypical] demonstrated that papers built on atypical combinations of prior work are more likely to be highly influential. This finding is based on a novelty index, which quantifies the degree of atypicality in the combinations of references at the journal level. More specifically, this index is calculated by comparing the observed co-occurrence of journal pairs in actual citation networks with their randomized co-occurrence under randomized citation networks. This comparison allows for computing the z-score of each possible journal pair cited by a given focal paper. A high and positive z-score indicates a conventional combination of journal pairs, while a low and negative z-score reflects a novel combination that deviates significantly from randomized co-occurrence patterns. Following the way that novelty is computed in the seminal work of Uzzi et al. [uzzi2013atypical] and SciSciNet  [lin2023sciscinet] dateset, we calculate the 10th percentile of the z-score distribution of the journal pairs cited by the focal paper, and apply its negative value to represent the 90th percentile of novelty. This approach underscores the relatively top atypical combinations of references within the paper, capturing its degree of novelty.
Interdisciplinarity
To assess the impact of a paper’s interdisciplinary nature, we apply a interdisciplinarity index for each paper. This index is calculated as the Rao–Stirling indicator  [stirling2007general, rafols2010diversity, zhang2016diversity, zhang2021scientometric, xiang2025evaluating] by combining three components: variety, balance, and disparity among the categories (e.g., scientific fields) to which a focal paper’s bibliography belongs. Specifically, for each field $i$, compute its share $p_{i}$ in the bibliography with whole counting method, which captures variety and balance. Then, for each pair of fields ($i,j$), calculate a distance measure $d_{ij}$ based on its co-citation frequency annually, which represents how intellectually far apart the fields are, capturing disparity in each year. Finally, the Rao–Stirling interdisciplinarity index is computed as:
$RS$ $=\sum\nolimits_{i,j(i\neq j)}^{n}p_{i}p_{j}d_{ij}.$ (2)
This formula increases when: more categories are represented, the distribution is more even, and the fields are cognitively more distant (less being co-cited).
@@
Lin et al.  [lin2023remote] reported the rise of remote collaborations that span locations and time zones. These collaborations are less likely to produce disruptive works as they tend to focus more on technical tasks rather than conceptual ones. The original calculation of team distance relies on the latitude and longitude of the affiliations listed in a paper’s byline. We follow this methodology to calculate the geographic team distance by the coordinate information of affiliations, which can further categorize teams into two types—onsite ($\leq 100$ km) and remote (>100>100 km) teams as defined by Lin et al  [lin2023remote].
@@
Zeng et al.  [zeng2021fresh] revealed a positive correlation between team freshness and scientific disruption, emphasizing the importance of new team dynamics in fostering innovative research. To quantify team freshness, we analyze the prior collaboration network of the authors involved in a given focal paper. In this network, the undirected edges represent prior collaborations between pairs of authors before coauthoring the focal paper. We classify team freshness into mutually exclusive categories based on the topological structure of the collaboration network, where $V$, $k$, and $n$ represent the node set, node degree, and network size, respectively. More specifically, the classification is as follows:
@@
  • 0: All authors have previously collaborated with each other, $\{\forall v_{i}\in V,\ k_{i}=n-1\}$;
  • 11: All authors have at least one prior collaboration link, $\{\forall v_{i}\in V,1\leq\ k_{i}\leq n-1\}$;
@@
  • 22: Some authors have at least one prior collaboration link, $\{\exists v_{i}\in V,\ k_{i}=0\}$;
  • 33: No author has any prior collaboration link, $\{\forall v_{i}\in V,\ k_{i}=0\}$.
Coarsened exact matching (CEM)
Coarsened Exact Matching (CEM) is a non-parametric matching method that directly coarsens the covariates into bins and then performs exact matching within these coarsened bins. This ensures excellent covariate balance within the matched samples by guaranteeing that all matched units have identical values on the coarsened covariates. The methodological procedure follows:
  • Treatment and Control Groups: We categorize papers into treated and control groups based on their KI values.
  • Covariates: We utilize the set of ex-ante covariates: focal paper properties (discipline and publication year), reference properties (novelty, interdisciplinarity, reference count, and average reference age), and team properties (size, geographic distance, and collaboration freshness).
  • Coarsening and matching: For categorical and numerical covariates with limited distinct values, like discipline and team freshness, we directly use their natural strata for binning. For other numerical covariates, we choose appropriate coarsening schemes by comparing the weighted $L_{1}$ imbalance and standardized mean difference (SMD); comprehensive analyses of coarsening schemes are displayed in Section S3. CEM then identified strata where exact matches on these coarsened covariates could be found. Only units within these matched strata were included in the subsequent analysis.
  • Weighting: To equalize the distribution of covariates between the groups, CEM weights are generated. Matched treated units are assigned a weight of $1$. Matched control units are weighted such that their distribution across strata mirrors that of the treated group. Formally, the weight for a control unit $i$ in stratum $s$ is defined as:
    $w_{i}=\frac{m_{C}}{m_{T}}\cdot\frac{m_{T}^{s}}{m_{C}^{s}}$
    where $m_{C}$ and $m_{T}$ are the total matched sample sizes in controlled and treated group, and $m_{C}^{s}$ and $m_{T}^{s}$ are the number of controlled and treated units in stratum $s$.
  • Effect estimation: To estimate the Average Treatment Effect on the Treated (ATT), we fit a Weighted Least Squares (WLS) regression model on the matched sample using the CEM-generated weights, which is specified as follows:
    $D=b+a_{ATT}\cdot T+\delta\cdot\rm Covariates+\epsilon$
    Where: $D$ is the outcome variable of disruption. $T$ is the binary treatment indicator of KI. Regarding doubly robust estimation, we control for the original Covariates to exploit within-stratum variation. The regression is weighted by the CEM weights ($w_{i}$). The coefficient of interest, $a_{ATT}$, captures the marginal effect of the treatment on the outcome after balancing ex-ante characteristics.
@@
To analyze the relationships between scientific disruption and various ex-ante covariates, we employ normalized ordinary least squares (OLS) regression models. This approach allows us to evaluate whether KI has a dominant effect on disruption. The regression models include the following covariates: focal paper properties (discipline and publication year), reference properties (novelty, interdisciplinarity, reference count, and average age), as well as team properties (size, geographic distance, and collaboration freshness). All numerical variables are standardized by z-score normalization to ensure comparability of effect coefficients across models. We start with a model that does not control for any covariates other than fixing discipline and publication year (3). Then, we control for each of the following covariates separately (4): focal paper’s novelty, interdisciplinarity, reference count, reference age, team size, team distance, and team freshness. Finally, we consider a model that accounts for all those covariates (5). The results are displayed in Table 1.
$D$ $=b_{0}+a_{0}\text{KI}+c_{0}F,$ (3)
$D$ $=b_{0}+a_{0}\text{KI}+a_{i}C_{i}+c_{0}F,$ (4)
$D$ $=b_{0}+a_{0}\text{KI}+\sum\nolimits_{i=1}^{7}a_{i}C_{i}+c_{0}F.$ (5)
To further quantify the relative importance of individual predictors within the full model (Model 9 in Table 1) and to systematically decompose the explained variance ($R^{2}$), we performed a dominance analysis  [azen2003dominance]. This method systematically assesses each predictor’s marginal contribution to $R^{2}$ across all possible sub-models, thereby attributing its unique and shared contributions to the total $R^{2}$. It establishes three types of dominance: conditional dominance (contribution in specific subsets), partial dominance (average contribution across specific subset sizes), and general dominance (overall average contribution across all subsets). The general dominance value, which inherently accounts for shared variance and avoids order dependence, serves as our primary measure of relative importance for each predictor. This approach allows for a robust and comprehensive comparison of the relative influence of all independent variables in the model.
Confounding Analysis
Confounding is a critical challenge in causal inference, where an ex-ante variable $C$ can distort the true effect of a target exposure $X$ on an outcome $Y$ ($C\rightarrow X,C\rightarrow Y$), leading to spurious associations or biased estimates ($X\dashrightarrow Y$). This analysis aims to rigorously estimate the causal effect of a target variable $X$ on an outcome $Y$ by identifying and adjusting for potential confounding factors that are common causes of both $X$ and $Y$, and are not causally affected by $X$. Here, we employ a model-based confounding analysis framework, comparing two parametric regression models. The first model, the Unadjusted Model (6), formulates the baseline association between the outcome $Y$ and the target variable $X$, while controlling for other relevant covariates $F$. The coefficient $\alpha_{1}$ in this model represents the unadjusted (or crude) effect of $X$ on $Y$. The second model, the Adjusted Model (7), expands upon the first by explicitly incorporating the identified confounder $C$ alongside $X$ and other covariates $F$. The coefficient $\delta_{1}$ in this model captures the adjusted effect of $X$ on $Y$, after accounting for the confounding effect of $C$. The presence and magnitude of confounding are primarily assessed by observing the change in the estimated coefficient of $X$ when the confounder $C$ is introduced into the model. The relative change in estimate (RCE) is defined as $|\frac{\alpha_{1}-\delta_{1}}{\alpha_{1}}|$. A substantial RCE indicates that $C$ indeed played a confounding role, distorting the unadjusted association between $X$ and $Y$. This adjustment helps to isolate $X$’s independent contribution to $Y$. In all conducted confounding analyses (i.e., citation impact as the target variable, and each of the ex-ante covariates as the potential confounder), we standardize the numerical variables by z-score normalization to ensure comparability of effect coefficients across models, with discipline and publication year controlled.
$Y$ $=\alpha_{0}+\alpha_{1}X+\alpha_{2}F,$ (6)
$Y$ $=\delta_{0}+\delta_{1}X+\delta_{2}C+\delta_{3}F.$ (7)
@@
Over the past decades, mediation analysis has seen rapid methodological development  [nguyen2021clarifying], becoming particularly prevalent in biomedical and social sciences. This approach is often applied to data from randomized controlled trials, aiming to investigate whether and how the effect of a target variable $X$ on an outcome $Y$ operates through a mediating variable $M$ ($X\rightarrow M\rightarrow Y$), where $M$ is an ex-post variable relative to $X$ and an ex-ante variable relative to $Y$. Here we utilize the model-based mediation framework, comprising two parametric regression models: Outcome Model (8) formulates the relation between the outcome $Y$ and the target variable $X$, as well as the potential mediator $M$, with covariates $F$ controlled. On the other hand, the Mediator Model (9) formulates the relation between the mediator $M$ and the target variable $X$, also controlling for covariates $F$. The indirect effect (ACME) is defined as the product of the coefficient $\beta_{1}$ of $X$ in model (9) and the coefficient $\theta_{2}$ of $M$ in model (8). This represents the portion of $X$’s effect on $Y$ mediated through $M$. The direct effect (ADE) is defined as the coefficient $\theta_{1}$ of $X$ in model (8), capturing the direct relationship between $X$ and $Y$, independent of $M$. The total effect (TE) is defined as the sum of the indirect effect (ACME) and direct effect (ADE). In all conducted mediation analyses (i.e., team composition as the target variable, and each of the ex-ante covariates as the potential mediator), we standardize the numerical variables by z-score normalization to ensure comparability of effect coefficients across models, with discipline and publication year controlled.
$Y$ $=\theta_{0}+\theta_{1}X+\theta_{2}M+\theta_{3}F,$ (8)
$M$ $=\beta_{0}+\beta_{1}X+\beta_{2}F.$ (9)
Acknowledgements
Funding:
National Natural Science Foundation of China (NSFC), No.72374173 (T. J.)
University Innovation Research Group of Chongqing, No.CXQT21005 (T. J.)
Fundamental Research Funds for the Central Universities, No.SWU-XDJH202303 (T. J.)
China Scholarship Council (CSC) program, No.202306990091 (X.-Y. Y.)
Chongqing Graduate Research Innovation Project, No.CYB22129 (X.-Y. Y.)
Author contributions:
Conceptualization: X.-Y. Y., T. R. and T. J.
Methodology: X.-Y. Y. and T. J.
Investigation: X.-Y. Y.
Visualization: X.-Y. Y.
Supervision: T. R. and T. J.
Writing—original draft, review & editing: X.-Y. Y., T. R. and T. J.
Competing interests:
The authors declare no competing interests.
Data and materials availability:
Correspondence and requests for materials should be addressed to T. R. or T. J.
Data availability: Our study draws on data from three publication datasets (OpenAlex  [priem2022openalex], SciSciNet  [lin2023sciscinet], and Web of Science) and one patent dataset (OECD-Patent). For OpenAlex dataset, you can access directly from https://docs.openalex.org/. For SciSciNet dataset [^1] We use version 1 of SciSciNet, as version 2 updates its publication records based on OpenAlex, the same source as our main dataset, which would result in data redundancy., you can access directly from https://doi.org/10.6084/m9.figshare.c.6076908.v1. The Web of Science dataset is not publicly available, and was used under license from their respective publishers. If you are interested, you may request access to the API through Clarivate, which requires an additional subscription or permission (https://clarivate.com/academia-government/scientific-and-academic-research/research-discovery-and-
referencing/web-of-science/web-of-science-core-collection/). For OECD-Patent dataset, you can access directly from https://www.oecd.org/en/data/datasets/intellectual-property-statistics.html.
Code availability: The computational codes for data processing, analysis, and visualization are available on a public repository (https://github.com/XiaoyaoYu95/Knowledge-Independence-Disruption).
Supplementary Materials
This PDF file includes:
Supplementary Text S1 to S8
Figures S1 to S31
Tables S1 to S17
@@ Acknowledgements
Figure 1: $|$ Distribution of reference count. The distribution of reference count for 87,419,190 articles recorded in OpenAlex follows a stretched exponential pattern, with a large number of papers containing relatively short reference lists.
@@
Figure 2: $|$ The association between disruption and KI is robust across different reference counts. (A-B) When controlling for the percentile of reference count, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference count. (C-D) When controlling for the reference count within the bottom group (<=10<=10), where each reference count induces a limited range of KI values, both the percentile and positive ratio of disruption continue to increase with KI across all groups of reference count.
@@
Figure 3: $|$ The association between disruption and KI is robust across alternative measures. (A-D) Upper panel: from the perspective of percentile, all disruption measures (Methods) show a consistent increasing trend with all KI measures. Lower panel: from the perspective of positive ratio, all disruption measures similarly increase with all KI measures. The positive ratio represents the likelihood of having a higher proportion of indind-type references compared to depdep-type references. Consequently, the variation curves for $D_{1}$ and $D_{3}$ align with those for $D_{0}$ and $D_{2}$, respectively.
@@
Figure 4: $|$ The association between disruption and KI is robust when controlling for novelty. (A) The percentile of novelty index (Methods) within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on less novel combinations of references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for more novel combinations of references. This pattern remains consistent across time periods. (B-C) When controlling for focal papers’ novelty, both the percentile and positive ratio of disruption continue to increase with KI across all levels of novelty.
@@
Figure 5: $|$ The association between disruption and KI is robust when controlling for novelty. (A) The percentile of interdisciplinarity index (Methods) within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on slightly less interdisciplinary combinations of references, which remains consistent across time periods. (B-C) When controlling for focal papers’ interdisciplinarity, both the percentile and positive ratio of disruption continue to increase with KI across all levels of interdisciplinarity.
@@
Figure 6: $|$ The association between disruption and KI is robust when controlling for reference age. (A) We represent the reference age of a given focal paper by calculating the the average age of cited references. The percentile of reference age within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers utilizing dependent or independent knowledge tend to rely on relatively younger references, whereas neutral papers—those maintaining a balance between dependent and independent knowledge—show a preference for older references. This pattern remains consistent across time periods. (B-C) When controlling for reference age, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference age.
Figure 7: $|$ The association between disruption and KI is robust when controlling for reference impact. (A) We represent the reference impact of a given focal paper by calculating the average $C_{5}$ of cited references. The percentile of reference impact within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite less impactful references. This pattern remains consistent across time periods. (B-C) When controlling for reference impact, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference impact. Additionally, papers citing less impactful references tend to exhibit slightly higher disruption.
Figure 8: $|$ The association between disruption and KI is robust when controlling for reference disruption. (A) We represent the reference disruption of a given focal paper by calculating the average disruption of cited references. The percentile of reference disruption within each bin is re-scaled by the average value of the corresponding period to highlight the trends. Papers that draw on more independent knowledge tend to cite more disruptive references. This pattern remains consistent across time periods. (B-C) When controlling for reference disruption, both the percentile and positive ratio of disruption continue to increase with KI across all levels of reference disruption. Additionally, papers citing more disruptive references tend to exhibit higher disruption.
@@
Figure 9: $|$ The association between disruption and KI is robust when controlling for team size. (A-B) When controlling for team size, both the percentile and positive ratio of disruption continue to increase with KI across different team sizes. Additionally, papers produced by smaller teams tend to exhibit slightly higher disruption.
@@
Figure 10: $|$ The association between disruption and KI is robust when controlling for team distance. (A-B) We calculate the geographic team distance by the coordinate information of affiliations (Methods). When controlling for team distance, both the percentile and positive ratio of disruption continue to increase with KI across different team distances. Additionally, papers produced by geographically closer teams tend to exhibit slightly higher disruption.
@@
Figure 11: $|$ The association between disruption and KI is robust when controlling for team freshness. (A-B) We classify the team freshness of a given focal paper by the topological structure of the prior collaboration network, where the value from 0 to 33 refers to network topology from a full-connected network to an isolated network (Methods). When controlling for team freshness, both the percentile and positive ratio of disruption continue to increase with KI across different team freshnesses. Additionally, papers produced by fresher teams tend to exhibit slightly higher disruption.